OpenStudy (anonymous):

Please Help x = y 2 - 1 2x - y = 4 Which of the following equations could be the result of using substitution to solve the system? 2y 2 - y - 5 = 0 2y 2 - y - 6 = 0 2y 2 - 2y - 5 = 0

OpenStudy (anonymous):

when using substation you should start with 2(y2-1) - y =4

OpenStudy (calculusxy):

\[\huge \tt \color{blue}{x = y^2 - 1}\] \[\huge \tt \color{blue}{2x - y = 4}\] you know the value of x as \(\large \tt \color{blue}{x = y^2 -1}\) so replace the x in the second equation with that to get: \[\huge \tt \color{green}{2(y^2 - 1) - y = 4}\]

OpenStudy (calculusxy):

Use the distributive property and add like terms, if necessary.

OpenStudy (calculusxy):

\[\huge \tt \color{red}{2y^2 - 2 -y = 4}\]

OpenStudy (calculusxy):

now move the 4 to the left side by doing the opposite (subtraction property of equality) \[\huge \tt \color{red}{2y^2 - 2 -y = 4} \rightarrow \huge \tt \color{green}{2y^2 - 6 - y = 0}\]

OpenStudy (calculusxy):

They put this as a quadratic: \[\huge \tt \color{purple}{ax^2 + bx + c= 0}\] We would rearrange this to put it into this format: \[\huge \tt \color{red}{2y^2 - y - 6 = 0}\]

OpenStudy (calculusxy):

@Austin1617